OpenStudy (pokemon23):

Four less than five times a number gives the same result as 6 added to three times the number. Find the number. Four less than five times a number gives the same result as 6 added to three times the number. Find the number. @Mathematics

OpenStudy (pokemon23):

I got 4-5x=6+3x

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

5x-4=3x=6

OpenStudy (anonymous):

5x -4 = 6+3x

OpenStudy (anonymous):

5x-4=3x+6 2x=10 x=5

OpenStudy (pokemon23):

who's correct?

OpenStudy (pokemon23):

oh ok thank you bath guy dude